trying to graph and label these systems of linear equations. State whether they are consistent, inconsistent dependent. If they are consistent, give the common solution.
2x-y=2
3x+y=18

Answer by ewatrrr(24785)

2x-y=2 P(0,-2) and (1,0) on this Line. Plot and connect with a line (Green)
3x+y=18 P(0,18) and (6,0) on this Line. Plot and connect with a line (Blue)
5x = 20
x = 4 and y = 6 8-y+=+2
P(4,6) the solution for this System
...........